Five years ago my age was four times as old as sheela .I am 25 at present what's the age of sheela
Answer by lwsshak3(11628)

Five years ago my age was four times as old as sheela .I am 25 at present what's the age of sheela
***
let x=present age of sheela
5 yrs ago:
x-5=sheela's age
25-5=my age
..
25-5=4(x-5)
4x-20=20
4x=40
x=10
present age of sheela=10